# A Weighted Hankel Approach and Cramér-Rao Bound Analysis for Quantitative Acoustic Microscopy Imaging

### v1

@@ Abstract
Quantitative acoustic microscopy (QAM) is a cutting-edge imaging modality that leverages very high-frequency ultrasound to characterize the acoustic and mechanical properties of biological tissues at microscopic resolutions. Radio-frequency echo signals are digitized and processed to yield two-dimensional maps. This paper introduces a weighted Hankel-based spectral method with a reweighting strategy to enhance robustness with regard to noise and reduce unreliable acoustic parameter estimates. Additionally, we derive, for the first time in QAM, Cramér-Rao bounds to establish theoretical performance benchmarks for acoustic parameter estimation. Simulations and experimental results demonstrate that the proposed method consistently outperform standard autoregressive approach, particularly under challenging conditions. These advancements promise to improve the accuracy and reliability of tissue characterization, enhancing the potential of QAM for biomedical applications.
{IEEEkeywords}
Quantitative acoustic microscopy, Hankel spectral method, Cramér-Rao bound, biomedical imaging
@@ 1 Introduction
[…]
Quantitative acoustic microscopy (QAM) employs ultrasound waves with frequencies exceeding 100 MHz to scan thin (5-14 $\mu$m thick) ex vivo tissue sections affixed to microscopy slides, using filtered water as the coupling medium [1]. This advanced imaging technique generates two-dimensional (2D) maps of acoustic properties, including speed of sound, impedance, and acoustic attenuation, in addition to mechanical properties such as bulk modulus, density, and compressibility, achieving spatial resolutions near the acoustic wavelength (e.g., finer than 8 $\mu$m at 250 MHz [2]). QAM offers unique and complementary advantages compared to conventional microscopy modalities (e.g., optical microscopy, electron microscopy, and histological imaging) for detailed tissue characterization [3]. Its versatility has been demonstrated in a wide range of applications, including corneal imaging [4], retinal studies [5, 6], liver tissue analysis [7], cerebellar folding [8], biomechanical changes in myopic eyes [9, 10], cancerous lymph nodes investigations [11], and living cells analysis [3, 12, 13].
Data acquired by QAM instruments consist of radio frequency (RF) signals containing two main reflections originating from the water-sample and sample-glass interfaces. These RF signals are then processed offline using model-based approaches to estimate acoustic parameters at each scanned position. While time-domain methods [12] have been explored, frequency-domain techniques have proven advantageous, particularly when reflections overlap in time [14]. The state-of-the-art in QAM, introduced by Rohrbach et al. [15], employs an autoregressive (AR) frequency-domain model with a spectral estimation algorithm inspired by the classical Prony method [16] and the annihilating filter [17].
@@
The first contribution, presented in Section 3, is the introduction of a weighted Hankel-based method that builds on Hankel matrix theory, as inspired by [18], to address the spectral estimation problem. In contrast to our previous approach [19], an iterative reweighting strategy is incorporated, to further improve robustness by mitigating the influence of outliers in the spectral data and minimize unreliable acoustic parameter estimates.
The second contribution, detailed in Section 4, is the derivation of the Cramér-Rao bounds presented for the first time in the context of QAM. These bounds establish theoretical performance benchmarks for evaluating and comparing the different estimators.
The third contribution is the rigorous evaluation of the proposed methods against the standard AR model through simulations and experimental data obtained from soft tissues using a QAM instrument operating at 500 MHz, as detailed in Section 5. The performance results, presented in Section 6, demonstrate that the proposed methodology achieves a significant reduction in unreliable estimates and improved estimation accuracy, all while incurring only a moderate additional computational cost. Finally, Section 7 discusses the findings and outlines potential directions for future research.
@@ 2 Signal Model in QAM
In QAM, an RF signal acquired in a tissue-free location results from a single reflection from the glass-water interface and serves as a reference signal, denoted by $h_{0}(t)$, where $t$ is the time (in seconds). On the other hand, a signal acquired at a tissue location, denoted by $h(t)$, is composed of two main reflections: one from the water-tissue interface ($h_{1}(t)$) and the other from the tissue-glass interface ($h_{2}(t)$). Fig. 1 illustrates the measurement principle in a tissue location (i.e., $h_{1}$ and $h_{2}$) and a tissue-free location (i.e., $h_{0}$).
In this context, the estimation of the acoustic properties of the sample relies on the accurate identification and characterization of $h_{1}$ and $h_{2}$ in terms of their amplitude, frequency attenuation, and time shift relative to $h_{0}$. This section provides an overview of the signal model, the estimation problem, and the image formation process involved in QAM technology.
@@ 2.1 From time to frequency domain
To account for the complex interactions of the acoustic waves within the tissue, including potential scattering effects and multiple reflections, we consider
@@
$h_{p}(t)=A_{p}h^{\beta_{p}}_{0}(t-t_{p}),$ (2)
where $A_{p}$ is the amplitude, $t_{p}$ is the time of flight (TOF) and ${\beta_{p}}$ symbolizes an additional frequency-dependent attenuation effect due to the round-trip propagation inside the sample.
@@
Let $\mathcal{F}_{h_{0}}$ and $\mathcal{F}_{h}$ denote the discrete Fourier transform (DFT) of the digital version of $h_{0}$ and $h$, respectively. The ratio between these two transforms at discrete frequencies $k\Delta f$ can be written as [15]
@@
where $\Delta f\!=\!f_{s}/(2M)$ denotes the frequency step size, $M$ represents the number of temporal samples (considering no zero-padding), and $f_{s}$ is the sampling frequency. In (3), the attenuation coefficient ${\beta_{p}}$ describes how the amplitude of the signal decreases with frequency, and the relative time shift of the $p$th echo compared to the reference signal is given by $\Delta t_{p}\!=\!t_{p}-t_{0}$.
Figure 1: Diagram of the experimental QAM approach. The measured signal $h$ comprises two components: $h_{1}$ and $h_{2}$, corresponding to reflections from the water-tissue and tissue-glass interfaces, respectively. The reference signal $h_{0}$ results from a single reflection at the glass-water interface in a tissue-free location.
@@ 2.2 Acoustic parameter estimation and image formation
Given $\{\lambda_{p}\}_{p=1}^{P}$ and $\{a_{p}\}_{p=1}^{P}$, the two leading reflections can be determined by reconstructing the $P$ pulses and computing the amplitude of the maximum of the Hilbert-transformed signals. The two pulses with the largest amplitudes correspond to the desired components related to the water-tissue and tissue-glass interfaces, which are denoted with the subscripts $p\!=\!1$ and $p\!=\!2$, without loss of generality. Then, the signal corresponding to water-sample interface is identified as the one with the smallest TOF ($t_{1}\!<\!t_{2}$), as it occurs before the sample-glass interface.
@@ 3 Weighted Hankel-based parameter estimation
The accurate estimation of $c$, $Z$, $\alpha$, and $d$ fundamentally depends on the precise determination of the spectral parameters $(\lambda_{1},a_{1})$ and $(\lambda_{2},a_{2})$. Although recovering amplitudes $a_{p}$ can be efficiently addressed using a least-squares (LS) approach, the primary challenge lies in accurately estimating $\lambda_{p}$.
To overcome the limitations of the standard AR-based framework for QAM [15], we previously introduced a Hankel (HK)-based spectral estimation framework for QAM image formation [19]. This method builds upon the HK spectral estimation technique originally presented in [18]. In this work, we extend our previous approach and propose a weighted HK-based method that incorporates a reweighting strategy, designed to better handle noisy data points within the normalized spectrum (i.e., (3)).
3.1 HK-based problem formulation
The HK spectral method applies to the parametric model (4) and seeks to approximate the vector $\boldsymbol{x}\!=\!(x_{\kappa_{1}},\ldots,x_{\kappa_{2N+1}})$ by another vector $\boldsymbol{g}\!=\!(g_{\kappa_{1}},\ldots,g_{\kappa_{2N+1}})$, whose elements are expressed as a linear combination of complex exponentials
$g_{\kappa_{n}}=\sum_{p=1}^{P}a_{p}\lambda_{p}^{\kappa_{n}}.$ (7)
@@ 3.1 HK-based problem formulation
A Hankel matrix $\boldsymbol{H}\!=\!H(\boldsymbol{g})$, generated element-wise from a vector $\boldsymbol{g}$, is a symmetric matrix where each ascending skew diagonal from left to right is constant such that $[\boldsymbol{H}]_{ij}\!=\!g_{\kappa_{i+j-1}}$ with $11\leq i,j\leq N+$ 1.
Kronecker’s theorem [21, 22] for complex symmetric matrices establishes that $\operatorname{rank}(\boldsymbol{H})\!=\!P$ if and only if there exist $\{\lambda_{p}\}_{p=1}^{P}$ and $\{a_{p}\}_{p=1}^{P}$ in $\mathbb{C}$ such that $g_{\kappa}\!=\!\sum_{p=1}^{P}a_{p}\lambda_{p}^{\kappa}$ for equally spaced indices $\kappa=\kappa_{1},\ldots,\kappa_{2N+1}$.
Consequently, the best approximation (in the $l_{2}$-norm sense) of $\boldsymbol{x}$ by a linear combination of $P$ complex exponentials is given by the vector $\boldsymbol{g}$ which satisfies $\operatorname{rank}(H(\boldsymbol{g}))=P$ and minimizes the $l_{2}$-norm of the difference $\boldsymbol{x}-\boldsymbol{g}$. Inspired by [18], $\boldsymbol{g}$ is obtained as the solution of the optimization problem
$\underset{{\boldsymbol{A},\boldsymbol{g}}}{\operatorname{minimize}}\quad\frac{1}{2}\boldsymbol{w^{T}}||\boldsymbol{x}-\boldsymbol{g}||_{2}^{2}+\mathcal{R}_{P}(\boldsymbol{H}),$ (8)
$\operatorname{subject\,to}\,\boldsymbol{H}=H(\boldsymbol{g}),$
where $\mathcal{R}_{P}(\boldsymbol{H})$ is an indicator function for square matrices, defined as $\mathcal{R}_{P}(\boldsymbol{H})=0$ if $\operatorname{rank}(\boldsymbol{H})\leq P$ and infinity otherwise, and ||.$||.||_{2}$ stands for $l_{2}$ vector norm. The weight vector $\boldsymbol{w}=(w_{\kappa_{1}},\ldots,w_{\kappa_{2N+1}})$ can be chosen depending on the specific application. In Section 3.4, we propose an iteratively reweighted approach to mitigate data noise.

### v2

@@ Abstract
Quantitative acoustic microscopy (QAM) is a cutting-edge imaging modality that leverages very high-frequency ultrasound to characterize the acoustic and mechanical properties of biological tissues at microscopic resolutions. Radio-frequency signals are digitized and processed to yield two-dimensional maps. This paper introduces a weighted Hankel-based spectral method with a reweighting strategy to enhance robustness with regard to noise and reduce unreliable acoustic parameter estimates. Additionally, we derive, for the first time in QAM, Cramér-Rao bounds to establish theoretical performance benchmarks for acoustic parameter estimation. Simulations and experimental results demonstrate that the proposed method consistently outperform standard autoregressive approach under challenging conditions. These advancements promise to improve the accuracy and reliability of tissue characterization, enhancing the potential of QAM for biomedical applications.
keywords:
Hankel spectral method , Cramér-Rao bound , Biomedical imaging , Quantitative acoustic microscopy
journal: Computers in Biology and Medicine
[…]
[label1]organization=Univ Lyon, INSA-Lyon, Université Claude Bernard Lyon 1, CNRS, Inserm, CREATIS UMR 5220, U1294, city=Villeurbanne, postcode=9100, country=France
[…]
[label2]organization=Department of Radiology,Weill Cornell Medicine, city=New York City, postcode=10022, state=NY, country=USA
[…]
[label3]organization=IRIT UMR CNRS 5505, Université Paul Sabatier Toulouse 3, city=Toulouse, postcode=31062, country=France
@@ 1 Introduction
Quantitative acoustic microscopy (QAM) employs ultrasound waves with frequencies exceeding 100 MHz to scan thin (5-14 $\mu$m thick) ex vivo tissue sections affixed to microscopy slides, using filtered water as the coupling medium [1, 2]. QAM is closely related to scanning acoustic microscopy (SAM), a well-established technique in nondestructive testing [3, 4] and in the study of hard biological samples (e.g., bone [5]). Although SAM has also been applied to soft tissue imaging [6, 7], providing insights into acoustic properties at frequencies near 100 MHz, its primary outputs are typically qualitative contrast maps, such as A-scans, B-scans, and C-scans.
In contrast to SAM, QAM produces two-dimensional (2D) quantitative maps of acoustic properties, including speed of sound, acoustic impedance, and attenuation, as well as mechanical properties such as bulk modulus, density, and compressibility, achieving spatial resolutions close to the acoustic wavelength (e.g., finer than 8 $\mu$m for a 250 MHz QAM system [8]). This transition from qualitative to quantitative imaging represents a significant advance, establishing QAM as a powerful and unique modality for biomedical research, offering complementary advantages over conventional microscopy modalities such as optical microscopy, electron microscopy, and histological imaging for detailed tissue characterization [9, 10]. Its versatility has been demonstrated in a wide range of applications, including corneal imaging [11], retinal studies [12, 8], liver tissue analysis [13], investigations of cerebellar folding [14], assessments of biomechanical changes in myopic eyes [15, 16], studies of cancerous lymph nodes [17], renal fibrosis [18], cutaneous melanoma [19], and analyses of living cells [9, 20, 21, 22]. Another study in the area highlights the relevance of acoustic parameters for tissue characterization and their dependence on spatial resolution [23].
Data acquired by QAM instruments consist of radio frequency (RF) signals containing two main reflections originating from the water-sample and sample-glass interfaces. These RF signals are then processed offline using model-based approaches to estimate acoustic parameters at each scanned position. While time domain methods [20] have been explored, frequency domain techniques have proven advantageous, particularly when the reflections overlap in time [24]. The current QAM methodology, introduced in [25], relies on an autoregressive (AR) frequency domain model combined with a spectral estimation algorithm inspired by the classical Prony method [26] and the annihilating filter [27].
@@
The first contribution, presented in Section 3, is the introduction of a weighted Hankel-based method that builds on Hankel matrix theory, as inspired by [28], to address the spectral estimation problem.
Related approaches that leverage structured low-rank modeling have been successfully applied to other imaging modalities, including computed tomography [29], spectroscopy [30], and image imputation tasks [31]. Structured low-rank methods have also been widely used in magnetic resonance imaging (MRI) reconstruction, where they have been employed to model spatially limited and phase-constrained images [32, 33], as well as transform-sparse images [34, 35]. For a comprehensive overview of structured low-rank modeling in MRI, see the review article [36].
Our approach shares several conceptual similarities with prior structured low-rank methods, although the specific matrix structures and application contexts differ. For instance, in MRI, k-space data are typically modeled using block-Hankel matrices, while in our case, a single Hankel matrix is constructed from the spectral data. Unlike in MRI where low-rank modeling is often central to image reconstruction, in QAM, spectral reconstruction via low-rank approximation primarily serves as an initial denoising step. Subsequent analysis requires identifying the two dominant signal components and estimating their parameters using an ESPRIT-type algorithm applied directly to the Hankel matrix formed from the denoised spectrum, rather than using a traditional covariance matrix, as is common in standard ESPRIT [37].
In contrast to fast MRI, where spectral data are frequently undersampled and require sophisticated reconstruction techniques, in QAM we work with a fully sampled spectrum. Additionally, our setting is inherently one-dimensional RF signals, whereas MRI typically involves two or three dimensional Fourier representations with geometric (gradient) information. These distinctions motivate our adaptations of existing structured low-rank approaches to suit QAM data.
A key contribution of this work is the introduction of a reweighting strategy designed to better manage noise-corrupted components in the normalized spectrum, thereby extending our earlier approach [38]. This is motivated by the trade-off between including a sufficient number of relevant frequency components and excluding corrupted ones, an issue that is particularly sensitive to the bandwidth selection.
While our approach focuses on reweighting the data fidelity term, it is worth noting that iterative reweighting strategies have also been proposed for regularizing the Hankel matrix itself, particularly in the context of low-rank priors in the Fourier domain for MRI reconstruction [39], and more generally in the context of low-rank matrix completion [40, 41]. However, to the best of our knowledge, our work represents the first adaptation of such techniques to the specific context and signal model of QAM data.
Another significant contribution of this work is the derivation of the Cramér-Rao bounds in the context of QAM, which is presented for the first time, as detailed in Section 4. These bounds set theoretical limits on the performance of estimators, enabling comparison of different estimation methods. They reveal the inherent challenges of estimating specific parameters under different conditions. They also facilitate the validation of physical modeling assumptions and help define experimental setups that promote reliable parameter estimation.
In addition, the proposed methods are compared against the standard AR model through extensive simulations and experimental data obtained from soft tissue using a QAM instrument operating at 500 MHz, as detailed in Section 5. The performance results, presented in Section 6, demonstrate that the proposed methodology significantly reduces unreliable estimates while maintaining or improving estimation accuracy under challenging conditions.
Finally, Section 7 discusses the findings and outlines potential directions for future research.
@@ 2 Signal Model in QAM
In QAM, an RF signal acquired in a tissue-free location results from a single reflection from the glass-water interface and serves as a reference signal, denoted by $h_{0}(t)$, where $t$ is the time (in seconds). On the other hand, a signal acquired at a tissue location, denoted by $h(t)$, is composed of two main reflections: one from the water-tissue interface $h_{1}(t)$ and the other from the tissue-glass interface $h_{2}(t)$. Fig. 1 illustrates the measurement principle in a tissue location (i.e., $h_{1}$ and $h_{2}$) and a tissue-free location (i.e., $h_{0}$).
In this context, the estimation of the acoustic properties of the sample relies on the accurate identification and characterization of $h_{1}(t)$ and $h_{2}(t)$ in terms of their amplitude, frequency attenuation, and time shift relative to  $h_{0}(t)$. This section provides an overview of the signal model, the estimation problem, and the image formation process involved in QAM technology.
(a) Experimental QAM approach
(b) Reference signal
(c) QAM signal
(d) Water-tissue echo
(e) Tissue-glass echo
Figure 1: (a) Diagram of the experimental QAM approach. The reference signal $h_{0}(t)$, shown in (b), is obtained from a single reflection at the glass-water interface in the absence of tissue. The measured signal $h(t)$ in (c) consists of two components, $h_{1}(t)$ and $h_{2}(t)$, associated with reflections at the water-tissue and tissue-glass interfaces, respectively. The echo signals $h_{1}(t)$ and $h_{2}(t)$ are shown in (d) and (e).
@@ 2.1 From time to frequency domain
We consider single layer samples, however to account for the complex interactions of the acoustic waves within the tissue, including potential scattering effects and multiple reflections, we consider
@@
$h_{p}(t)=A_{p}h^{{*}}_{0}(t-t_{p}),$ (2)
where $A_{p}$ is the real amplitude, $t_{p}$ is the time of flight (TOF) and * symbolizes an additional frequency-dependent attenuation effect due to the round-trip propagation inside the sample. For consistency in the notation, the signals related to the water-tissue and tissue-glass interfaces remain denoted by the subscripts $p\!=\!1$ and $p\!=\!2$.
@@
Taking the Fourier transform (FT) of (1)-(2), and assuming frequency-dependent attenuation is linear with both frequency and propagation distance, we obtain
$\mathcal{F}_{h}(f)=\operatorname{FT}(h(t))=\mathcal{F}_{h_{0}}(f)\sum_{p=1}^{P}{A_{p}}\exp{(2\pi f[-{\beta_{p}}-\mathrm{i}{\Delta t_{p}}])}.$ (3)
Here, the “*” were dropped to introduce the explicit frequency-dependent attenuation coefficients $\beta_{p}$ (in Np/MHz). It is also assumed that $\beta_{1}=0$ Np/MHz, implying negligible attenuation for the reflection from the water-tissue interface. In (3), the relative time shift of the $p$th echo compared to the reference signal is given by $\Delta t_{p}\!=\!t_{p}-t_{0}$.
The ratio between $\mathcal{F}_{h}$ and $\mathcal{F}_{h_{0}}$ at discrete frequencies $k\Delta f$ can be written as
@@
where $\Delta f\!=\!f_{s}/(2M)$ denotes the frequency step size, $M$ represents the number of points in the sampled signal (considering no zero-padding), and $f_{s}$ is the sampling rate.
@@ 2.2 Acoustic parameter estimation and image formation
Given $\{\lambda_{p}\}_{p=1}^{P}$ and $\{a_{p}\}_{p=1}^{P}$, the two leading reflections can be determined by reconstructing the $P$ pulses and computing the amplitude of the maximum of the Hilbert-transformed signals. The two pulses with the largest amplitudes correspond to the desired components related to the water-tissue and tissue-glass interfaces, which are denoted with the subscripts $p\!=\!1$ and $p\!=\!2$, without loss of generality. Then, the signal corresponding to water-sample interface is identified as the one with the smallest TOF ($t_{1}\!<\!t_{2}$), as it occurs before the sample-glass interface [25].
@@ 3 Weighted Hankel-based parameter estimation
The accurate estimation of $c$, $Z$, $\alpha$ and $d$ depends fundamentally on the precise determination of the spectral parameters $(\lambda_{1},a_{1})$ and $(\lambda_{2},a_{2})$. Once $\lambda_{p}$ is known, the amplitudes $a_{p}$ can be recovered efficiently using a least squares (LS) approach. However, accurate estimation of $\lambda_{p}$ can be challenging.
To overcome the limitations of the standard AR-based framework for QAM [25], we previously introduced a Hankel-based spectral estimation framework for QAM image formation [38]. This method builds upon the spectral estimation technique originally presented in [28]. In contrast to our previous approach [38], an iterative reweighting strategy is incorporated, to further improve robustness by mitigating the influence of outliers in the spectral data and minimize unreliable acoustic parameter estimates.
3.1 Hankel-based problem formulation
The Hankel-based spectral method applies to the parametric model (5) and seeks to approximate the vector $\boldsymbol{x}\!=\!(x_{\kappa_{1}},\ldots,x_{\kappa_{2N+1}})$ by $\boldsymbol{g}\!=\!(g_{\kappa_{1}},\ldots,g_{\kappa_{2N+1}})$, whose elements are expressed as a linear combination of complex exponentials
$g_{\kappa_{n}}=\sum_{p=1}^{P}a_{p}\lambda_{p}^{\kappa_{n}},$ (8)
@@ 3.1 Hankel-based problem formulation
Using Kronecker’s theorem for complex symmetric matrices, the solution to this approximation problem can be expressed in terms of $\boldsymbol{g}$, whose associated Hankel matrix has rank $P$ and minimizes the residual between $\boldsymbol{x}$ and $\boldsymbol{g}$.
For the sake of completeness, let’s recall that a Hankel matrix $\boldsymbol{H}\!=\!\mathcal{H}(\boldsymbol{g})$, generated element-wise from a vector $\boldsymbol{g}$, is a symmetric matrix where each ascending skew diagonal from left to right is constant such that $[\boldsymbol{H}]_{ij}\!=\!g_{\kappa_{i+j-1}}$ with $11\leq i,j\leq N+$ 1. On the other hand, Kronecker’s theorem [43, 44] for complex symmetric matrices establishes that $\operatorname{rank}(\boldsymbol{H})\!=\!P$ if and only if there exist $\{\lambda_{p}\}_{p=1}^{P}$ and $\{a_{p}\}_{p=1}^{P}$ in $\mathbb{C}$ such that $g_{\kappa}\!=\!\sum_{p=1}^{P}a_{p}\lambda_{p}^{\kappa}$ for equally spaced indices $\kappa=\kappa_{1},\ldots,\kappa_{2N+1}$.
Consequently, the best approximation (in the $l_{2}$-norm sense) of $\boldsymbol{x}$ by a linear combination of $P$ complex exponentials is given by the vector $\boldsymbol{g}$ which satisfies $\operatorname{rank}(\mathcal{H}(\boldsymbol{g}))=P$ and minimizes the $l_{2}$-norm of the difference $\boldsymbol{x}-\boldsymbol{g}$. Inspired by [28], $\boldsymbol{g}$ is obtained as the solution of the optimization problem
$\underset{\boldsymbol{H},\,\boldsymbol{g}}{\operatorname{minimize}}\quad{\frac{1}{2}\,\|\boldsymbol{x}-\boldsymbol{g}\|_{\boldsymbol{W}}^{2}}+\mathcal{R}_{P}(\boldsymbol{H}),$ (9)
$\operatorname{subject\,to}\quad\boldsymbol{H}=\mathcal{H}(\boldsymbol{g}),$
where $\mathcal{R}_{P}(\boldsymbol{H})$ is an indicator function that equals zero if $\operatorname{rank}(\boldsymbol{H})\leq P$, and infinity otherwise, $\|\boldsymbol{x}-\boldsymbol{g}\|_{\boldsymbol{W}}^{2}\!=\!(\boldsymbol{x}-\boldsymbol{g})^{\mathrm{H}}\boldsymbol{W}(\boldsymbol{x}-\boldsymbol{g})$ stands for the $\boldsymbol{W}$-weighted $l_{2}$-norm and $\cdot^{H}$ is the conjugate transpose. The matrix $\boldsymbol{W}\!=\!\operatorname{diag}(\boldsymbol{w})$ is a diagonal matrix with non-negative entries defined by the weight vector $\boldsymbol{w}=(w_{\kappa_{1}},\ldots,w_{\kappa_{2N+1}})$. In Section 3.4, we propose an iteratively reweighted approach to mitigate noisy data, in contrast to our previous approach introduced in [38], which assumes no weights, i.e., $\boldsymbol{w}\equiv\boldsymbol{1}$.